Find the equation of a line which has (-4,0) and (0,2) x and y intercept respestively
Answer by Alan3354(69443)

Find the equation of a line which has (-4,0) and (0,2) x and y intercept respestively
------------------
For intercepts (A,0) & (0,B), it's
Bx + Ay = A*B
2x - 4y = -8
x - 2y = -4